# Physics Prospects with MeV Neutrino-Argon Charged Current Interactions using Enhanced Photon Detection in Future LArTPCs

### v1

@@ I Introduction
Noble liquid time projection chambers (TPCs) are widely employed for detecting rare and low-energy signals, such as neutrinos and dark matter interactions. Liquid argon (LAr) is particularly favored for its scalability and excellent tracking and calorimetric capabilities [1]. Experiments such as LArIAT [2] demonstrate sensitivity to neutrinos with energies as low as a few MeV, where electron neutrino ($\nu_{e}$) charged current (CC) interactions on argon nuclei serve as the primary detection channel. These neutrinos originate from supernova bursts, solar fusion reactions, and the diffuse supernova neutrino background (DSNB) [3, 4, 5].
@@
Argon (Ar) and xenon (Xe) emit vacuum ultra-violet (VUV) scintillation light, peaking at 127 nm and 178 nm, respectively, when charged particles deposit energy inside [6, 7]. In LArTPCs, charge signals from ionization can be collected in the presence of an electric field. The anti-correlation between charge and light signals, well-established in liquid xenon (LXe) experiments, has been used to improve energy reconstruction in searches for neutrinoless double beta decay and dark matter detection [8, 9]. In the XENON experiment, the hypothesized weakly interacting massive particle dark matter candidate would be detected through nuclear or electron recoils that deposit tens of keV energy [10, 11]. EXO-200 [8] has performed measurements of the absolute yields of charge and light in LXe for electron recoils up to MeV.
@@
While charge detection in LArTPCs has matured [12, 13, 14, 15, 16], optical coverage and light yield remain limited. In LArTPC neutrino experiments, event reconstruction primarily uses the charge signal while the light signal is mostly used to provide timing information for non-beam neutrino events. Matching charge and light signals to facilitate cosmic background rejection has been explored at MicroBooNE [17]. It is anticipated an enhanced photon detection system in future LArTPCs can offer more capabilities for physics with MeV $\nu_{e}$-Ar CC interactions. The recently proposed Aluminum Profiles with Embedded X-Arapuca (APEX) [18] suggests optical coverage of up to 60% ($\sim$2000 m2) and an average light yield of $\sim$180 photoelectrons per MeV (PE/MeV), presenting a great potential for low-energy physics.
@@
In this work, we study the prospects of MeV neutrino physics in a LArTPC with enhanced photon detection. Key improvements include increased MeV energy resolution from light calorimetry, excellent timing resolution, and extensive optical coverage. These advancements facilitate the identification of delayed energy deposits and particle tagging, improving signal detection and background rejection, particularly in searches for the DSNB as well as many other interesting physics topics [19, 20, 21, 22, 23, 24, 25, 26]. The insights gained in this study will guide the design and calibration of photon detection systems in future LArTPCs.
@@ II Simulation
The MeV-scale $\nu_{e}$ CC interactions with an Ar nucleus are simulated using the MARLEY v1.2.0 event generator [27, 28]. Monoenergetic events are generated at energies ranging from 5 MeV to 50 MeV in 5 MeV increments. These events originate at the coordinate origin and propagate in the $\hat{z}$ direction. For each energy, 1000 events are generated. The final-state particles from the neutrino interaction are tracked using Geant4 v4.10.6.p01 [29, 30, 31], interfaced via the edep-sim package [32]. The simulation is conducted in a large LAr volume, measuring 200 meters in each dimension, to ensure full containment of all simulated events. The deposited energy is recorded with a maximum step size of 0.5 cm, along with the corresponding timing information for each energy deposit.
@@
The two-dimensional distribution of the true $\nu_{e}$ energy ($E_{\nu_{e}}$) versus $E_{\text{avail}}$ is shown in Fig. 2. Starting at 10 MeV, an energy loss of approximately 8 MeV due to nuclear binding energy is observed when hadrons are ejected from the Ar nucleus. This results in a secondary peak in the $E_{\text{avail}}$ distribution, offset by 8 MeV from the primary peak (cf. Fig. 3c). Above 20 MeV, multiple hadrons can be ejected from the nucleus. Overall, hadron knockout events constitute about 25% of all $\nu_{e}$ CC interactions across the 5-50 MeV range, with neutrons being the most common ($\sim$65%). The neutron’s 7.9 MeV binding energy escapes detection. Protons and $\alpha$ particles also appear in the final-state but are less common, requiring 7.6 MeV and 7.1 MeV, respectively, to overcome their binding energies. Additionally, a systematic energy shift of approximately --1 MeV is observed in the MARLEY event generator, accounting for the binding energy change from the initial Ar nucleus to the K nucleus [28]. Nuclear quenching effects on the small, keV-scale recoil energy of the daughter 40K nucleus are negligible and are therefore ignored in this study.
Figure 2: Distribution of $E_{\text{avail}}$ and true $E_{\nu_{e}}$ for generated $\nu_{e}$-Ar CC interactions for incoming $\nu_{e}$ with energy in 5 - 50 MeV.
@@
For each final-state particle interacting with the Ar nucleus, we define its deposited energy, $E_{\text{depo}}$, which includes energy deposition from all daughter particles involved in scattering, decay, de-excitation, etc. The primary energy smearing effect at this stage arises from neutron capture on the argon-40 (40Ar) nucleus, where an additional 6.1 MeV is deposited due to the de-excitation of the resulting argon-41 (41Ar) nucleus [34]. In this study, only neutron captures on 40Ar are simulated, and all neutrons in the simulated events are eventually captured. The 41Ar de-excitation releases several low energy gammas which may be separated out using timing and spatial information but is not considered here.
@@
The secondary peak in the $E_{\text{avail}}$ distribution (Fig. 3c) further splits into two smaller secondary peaks in the $E_{\text{depo}}$ distribution. Events containing neutrons are clustered approximately 1.8 MeV below the primary $E_{\text{depo}}$ peak due to the combined effects of undetected neutron binding energy and the additional energy released from neutron capture. Furthermore, the $E_{\text{depo}}$ of most events is shifted downward by 0.51 MeV relative to $E_{\text{avail}}$, corresponding to the rest mass of the final-state electron.
Figure 3: Deposited energy distribution for charge (a), light (b), and combined calorimetry (c) for 35 MeV $\nu_{e}$-Ar CC events.
@@
In summary, two energy smearing mechanisms are relevant for MeV-scale $\nu_{e}$ CC events. The first is the hadron knockout effect from the primary $\nu_{e}$-Ar interaction, where a binding energy of approximately 8 MeV remains undetected in higher energy $\nu_{e}$ events. The second arises from the neutron capture, which introduces an additional 6.1 MeV deposition due to the $\gamma$ cascade from the de-excitation of the 41Ar. These two mechanisms create secondary peaks at lower $E_{\text{depo}}$ (Fig. 3c), broadening the energy distribution and impacting low-energy physics searches.
@@ III Charge and Light Detection
Two charge detection thresholds, 75 keV and 500 keV, are considered and applied to the simulated charge calorimetry. The 75 keV threshold represents an optimistic detection limit, similarly used in Ref [33]. MicroBooNE has achieved a 210 keV charge detection threshold, where the reconstruction efficiency for low energy electrons reaches 50% of its maximum achievable value in a “low-threshold” configuration on its charge collection plane [42]. The 500 keV threshold serves as a more conservative benchmark. For each energy deposit dQ, if it falls below the selected threshold, it is excluded from the charge calorimetry of the event.
@@
where dQ comes from Eq. 1. The number of VUV photons is then simulated based on dL and the assumed light yields. We consider several benchmark detector average light yields ($\overline{\text{LY}}$) from 35 to 220 photoelectrons per MeV of deposited energy (PE/MeV). The lower bound of 35 PE/MeV is chosen because it is close to the average light yield of the LArTPC module under construction [43]. Higher $\overline{\text{LY}}$ values, up to 220 PE/MeV, are studied to explore what $\overline{\text{LY}}$ future LArTPC should aim to achieve. In LAr, approximately 21622 photons are produced per MeV of deposited energy for a minimum ionizing particle (MIP) with ${R}_{\text{c}}=$0.7 at a typical electric field of 0.5 kV/cm [35]. The overall photon collection efficiency (PCE) is defined as the ratio between the number of detected photoelectrons in the full photon detection system and the total number of photons initially produced. For each benchmark $\overline{\text{LY}}$, the corresponding PCE is calculated as:
@@
Table 1: The photon collection efficiency (PCE) of five assumed average light yields.
The simulated expected average number of produced photons, $\overline{{N}}_{\text{ph}}$, in LAr is:
$\overline{{N}}_{\text{ph}}=\frac{{dL}}{19.5\ \text{eV}}\times\text{PCE}$ (5)
@@
${N}_{\text{PE}}=\text{Poisson}(\overline{{N}}_{\text{ph}},\sqrt{\overline{{N}}_{\text{ph}}})$ (6)
The final detected energy deposited in light is then derived from the detected photoelectrons:
@@
Fig. 3 shows the distributions of deposited charge and light from 35 MeV $\nu_{e}$ CC interactions on Ar. In addition to ${E}_{\text{avail}}$ and ${E}_{\text{depo}}$ described in Sec. II, the energy deposited as charge at zero charge detection threshold, ${Q}_{0}$, primarily peaks around 19 MeV (Fig. 3a). Secondary peaks in the ${Q}_{0}$ distribution originate from ${E}_{\text{depo}}$ due to events with knockout hadrons. The secondary peak at 18 MeV in ${Q}_{0}$ corresponds to events with neutron emission, while the broader peak at 12 MeV consists of events with a proton or an $\alpha$ particle. Compared to ${E}_{\text{depo}}$, the Birks model introduces additional broadening in ${Q}_{0}$. This broadening is more pronounced for secondary peaks because knockout hadrons are not MIPs, leading to greater variation in recombination effects as described by the Birks model.
@@
For energy deposited as light, the events are centered around 14 MeV (Fig. 3b). The secondary peaks are less distinct and merge with the primary peak. This effect is further illustrated in Fig. 4, which shows the energy deposited in light for $\overline{\text{LY}}$ = 220 PE/MeV (${L}_{220}$). Specifically, for events with neutrons, the net bias in deposited energy is --1.8 MeV. The total deposited energy of these events typically fall within the two-standard-deviation band of the primary light peak. Consequently, neutron emission events populate the lower tail of the $\text{L}_{220}$ distribution in Fig. 4 rather than forming a distinct secondary peak, as observed in the charge distribution.
It is important to note that nuclear quenching effects in hadron-nucleus interactions are not explicitly simulated in this study. In particular, a lower scintillation efficiency has been observed in tens of keV nuclear recoils in LXe [44]. For $\alpha$ particles, a smaller light yield is observed possibly explained by a biexicitonic quenching process; while for few MeV protons, this was calculated to be negligible [45]. For neutrons in this study, the potential extra light loss could shift neutron emission events further from the primary peak in the light distribution.
The single peak structure of the light distribution is observed across all $\nu_{e}$ samples from 5 to 50 MeV and holds true for all benchmark $\overline{\text{LY}}$. Unlike the charge calorimetry, no systematic energy bias is introduced when varying $\overline{\text{LY}}$. However, energy smearing effects become more pronounced as $\overline{\text{LY}}$ decreases from 220 PE/MeV to 35 PE/MeV.
Finally, the event ${E}_{\text{depo}}$ is largely recovered when combining charge and light calorimetry, ${Q}_{0}$ + ${L}_{220}$ (Fig. 3c). Similar to charge-only calorimetry, the application of a 75 keV charge detection threshold introduces spectral broadening and an energy bias in the combined calorimetry. Additionally, implementing the modified Box model[46] to simulate recombination effects yields similar results to those obtained using the Birks model.
@@ IV.1 Energy Reconstruction
In this section, we examine the potential for energy reconstruction using charge and light calorimetry in a future LArTPC with an enhanced photon detection system. The ${E}_{\text{avail}}$ for $\nu_{e}$ CC events can be reconstructed by applying a scaling factor to the deposited energy. The scaling factor represents the fraction of ${E}_{\text{avail}}$ that is converted into deposited energy in charge, light, or combined calorimetry in each event.
As an example, we consider charge calorimetry with a detection threshold of 75 keV, ${Q}_{75}$. From simulation, we determine a scale factor of 0.54, taken as the mode of the ratio distribution between ${Q}_{75}$ and ${E}_{\text{avail}}$. The reconstructed energy from ${Q}_{75}$ is then given by:
@@
The energy distributions obtained using these three reconstruction methods for 35 MeV $\nu_{e}$-Ar CC events are shown in Fig. 5. Notably, the combined charge and light energy reconstruction (Eq. 10) results in a narrower distribution than charge-only reconstruction (Eq. 8). Although secondary peaks, inherited from the charge-only energy distribution, remain present, the primary peak in the combined calorimetry distribution is significantly narrower than that of the light-only reconstruction (Eq. 9).

### v2

@@ I Introduction
Noble liquid time projection chambers (TPCs) are widely employed for detecting rare and low-energy signals, such as neutrinos and dark matter interactions. Liquid argon (LAr) is particularly favored for its scalability and excellent tracking and calorimetric capabilities [1]. Experiments such as LArIAT, ArgoNeuT, and MicroBooNE [2, 3, 4] have demonstrated potentials of reconstructing MeV-scale activities. This capability is crucial for the detection and identification of neutrinos with energies as low as a few MeV, where electron neutrino ($\nu_{e}$) charged current (CC) interactions on argon nuclei serve as the primary detection channel at DUNE [5]. These neutrinos originate from supernova bursts [6, 7], solar fusion reactions [8, 9, 10], and the diffuse supernova neutrino background (DSNB) [11, 12, 13].
@@
Argon (Ar) and xenon (Xe) emit vacuum ultra-violet (VUV) scintillation light, peaking at 127 nm and 178 nm, respectively, when charged particles deposit energy inside [14, 15]. In noble liquid TPCs, charge signals from ionization can be collected in the presence of an electric field. The anti-correlation between charge and light signals, well-established in liquid xenon (LXe) experiments, has been used to improve energy reconstruction in searches for neutrinoless double beta decay and dark matter [16, 17]. In the XENON experiment, the hypothesized weakly interacting massive particle dark matter candidate would be detected through nuclear or electron recoils that deposit tens of keV energy [18, 19]. EXO-200 [16] has performed measurements of the absolute yields of charge and light in LXe for electron recoils up to MeV.
@@
While charge detection in LArTPCs has matured [20, 21, 22, 23, 24], optical coverage and light yield (LY) remain limited. In LArTPC neutrino experiments, event reconstruction primarily uses the charge signal while the light signal is mostly used to provide timing information for non-beam neutrino events. Matching charge and light signals to facilitate cosmic background rejection has been explored at MicroBooNE [25]. It is anticipated an enhanced photon detection system in future LArTPCs can offer more capabilities for physics with MeV $\nu_{e}$-Ar CC interactions. The recently proposed Aluminum Profiles with Embedded X-Arapuca (APEX) [26] for DUNE’s third and fourth far detector LArTPC modules suggests optical coverage of up to 60% ($\sim$2000 m2) and an average LY of $\sim$180 photoelectrons per MeV (PE/MeV), presenting a great potential for low-energy physics.
@@
In this work, we study the prospects of MeV neutrino physics in a LArTPC with enhanced photon detection. Key improvements include increased MeV energy resolution from light calorimetry, excellent timing resolution, and extensive optical coverage. These advancements facilitate the identification of delayed energy deposits and particle tagging, improving signal detection and background rejection, particularly in searches for the DSNB as well as many other interesting physics topics [9, 5]. The insights gained in this study will guide the design and calibration of photon detection systems in future LArTPCs.
@@ II Simulation
The MeV-scale $\nu_{e}$ CC interactions with an Ar nucleus are simulated using the MARLEY v1.2.0 event generator [27, 28]. Monoenergetic events are generated at energies ranging from 5 MeV to 50 MeV in 5 MeV increments. These events originate at the coordinate origin and propagate in the $\hat{z}$ direction. For each energy, 1000 events are generated. The final-state particles from the neutrino interaction are tracked using Geant4 v4.10.6.p01 [29, 30, 31], interfaced via the edep-sim package [32]. The default $\tt{QGSP\_BERT}$ physics list is used. The neutron killer is deactivated to properly simulate neutron and its captures. The simulation is conducted in a large LAr volume, measuring 200 meters in each dimension, to ensure full containment of all simulated events. The deposited energy is recorded with a maximum step size of 0.5 cm, along with the corresponding timing information for each energy deposit. All electrons from energy deposition into charge calorimetry are assumed to drift under an infinite lifetime and be collected by charge readout.
@@
The two-dimensional distribution of the true $\nu_{e}$ energy ($E_{\nu_{e}}$) versus $E_{\text{avail}}$ is shown in the top plot of Fig. 2. Starting at 10 MeV, an energy loss of approximately 8 MeV due to nuclear binding energy is observed when hadrons are ejected from the Ar nucleus. This results in a secondary peak in the $E_{\text{avail}}$ distribution, offset by 8 MeV from the primary peak (cf. Fig. 3 bottom plot). The fraction of hadron production events as a function of true $E_{\nu_{e}}$ is shown in the bottom plot of Fig. 2. All hadrons considered in the study below–neutrons, protons, and $\alpha$ particles–have higher production rates as neutrino energy increases. Overall, hadron knockout events constitute about 25% of all $\nu_{e}$ CC interactions across the 5-50 MeV range, with neutrons being the most common ($\sim$65%). The neutron’s 7.9 MeV binding energy escapes detection. Protons and $\alpha$ particles also appear in the final-state but are less common, requiring 7.6 MeV and 7.1 MeV, respectively, to overcome their binding energies. Additionally, a systematic energy shift of approximately --1 MeV is observed in the MARLEY event generator, accounting for the binding energy change from the initial Ar nucleus to the K nucleus [28]. Nuclear quenching effects on the small, keV-scale recoil energy of the daughter 40K nucleus are negligible and are therefore ignored in this study.
Figure 2: Top: Distribution of $E_{\text{avail}}$ and true $E_{\nu_{e}}$ for generated $\nu_{e}$-Ar CC interactions for incoming $\nu_{e}$ with energy in 5 - 50 MeV. Bottom: Fraction of hadron production events as a function of true neutrino energy. Most hadron emission events produce only a neutron (green, $\text{N}_{\text{n}}$ = 1), a proton (red, $\text{N}_{\text{p}}$ = 1), or an alpha particle (pink, $\text{N}_{\alpha}$ = 1). The black line shows the fraction of events with a production of any of the three hadrons.
@@
For each final-state particle exiting the nucleus and interacting with the bulk argon, we define its deposited energy, $E_{\text{depo}}$, which includes energy deposition from all daughter particles involved in scattering, decay, de-excitation, etc. The primary energy smearing effect at this stage arises from neutron capture on the argon-40 (40Ar) nucleus, where an additional 6.1 MeV is deposited due to the de-excitation of the resulting argon-41 (41Ar) nucleus [34]. In this study, only neutron captures on 40Ar are simulated, and all neutrons in the simulated events are eventually captured. The 41Ar de-excitation releases several low energy gammas which may be separated out using timing and spatial information but is not considered here.
@@
The secondary peak in the $E_{\text{avail}}$ distribution (Fig. 3 bottom plot) further splits into two smaller secondary peaks in the $E_{\text{depo}}$ distribution. Events containing neutrons are clustered approximately 1.8 MeV below the primary $E_{\text{depo}}$ peak due to the combined effects of undetected neutron binding energy and the additional energy released from neutron capture. Events containing hadrons other than neutron cluster at $E_{\text{depo}}$ regions similar to the secondary peak in the $E_{\text{avail}}$ distribution. Furthermore, the $E_{\text{depo}}$ of most events is shifted downward by 0.51 MeV relative to $E_{\text{avail}}$, corresponding to the rest mass of the final-state electron.
@@
In summary, two energy smearing mechanisms are relevant for MeV-scale $\nu_{e}$ CC events. The first is the hadron knockout effect from the primary $\nu_{e}$-Ar interaction, where a binding energy of approximately 8 MeV remains undetected in higher energy $\nu_{e}$ events. The second arises from the neutron capture, which introduces an additional 6.1 MeV deposition due to the $\gamma$ cascade from the de-excitation of the 41Ar. These two mechanisms create secondary peaks at lower $E_{\text{depo}}$ (Fig. 3 bottom plot), broadening the energy distribution and impacting low-energy physics searches.
@@ III Charge and Light Detection
Two charge detection thresholds, 75 keV and 500 keV, are considered and applied to the simulated charge calorimetry. The 75 keV threshold represents an optimistic detection limit, similarly used in Ref [33]. MicroBooNE has achieved a 210 keV charge detection threshold, where the reconstruction efficiency for low energy electrons reaches 50% of its maximum achievable value in a “low-threshold” configuration on its charge collection plane [42]. The 500 keV threshold serves as a more conservative benchmark. For each single energy deposit ($dE$) into charge calorimetry ($dQ$), if $dQ$ falls below the selected threshold, it is excluded from the charge calorimetry of the event.
@@
where dQ comes from Eq. 1. The number of VUV photons is then simulated based on dL and the assumed LY. We consider several benchmark detector average light yields ($\overline{\text{LY}}$) from 35 to 220 photoelectrons per MeV of deposited energy (PE/MeV). The lower bound of 35 PE/MeV is chosen because it is close to the ($\overline{\text{LY}}$) of the DUNE vertical drift far detector under construction [43]. Higher $\overline{\text{LY}}$ values, up to 220 PE/MeV, are studied to explore what $\overline{\text{LY}}$ future LArTPC should aim to achieve. In LAr, approximately 21622 photons are produced per MeV of deposited energy for a minimum ionizing particle (MIP) with ${R}_{\text{c}}=$0.7 at a typical electric field of 0.5 kV/cm [35]. The overall photon collection efficiency (PCE) is defined as the ratio between the number of detected photoelectrons in the full photon detection system and the total number of photons initially produced. For each benchmark $\overline{\text{LY}}$, the corresponding PCE is calculated as:
@@
Table 1: The photon collection efficiency (PCE) of five assumed $\overline{\text{LY}}$.
For each individual energy deposition, the simulated expected average number of produced photons, $\overline{{N}}_{\text{ph}}$, in LAr is:
$\overline{{N}}_{\text{ph}}=\frac{{dL}}{19.5\ \text{eV}}\times\text{PCE}(\overline{\text{LY}})$ (5)
@@
${N}_{\text{PE}}=\text{Poisson}(\overline{{N}}_{\text{ph}})$ (6)
The final detected energy deposited in light is then derived from the detected photoelectrons for each individual energy deposition:
@@
and summed across the event. A similar calculation taking into account the non-uniform LY and position reconstruction uncertainty is detailed in Appendix B.
Figure 3: Deposited energy distribution for charge (top), light (middle), and combined calorimetry (bottom) for 35 MeV $\nu_{e}$-Ar CC events. In the legend, the energy deposited as charge calorimetry at zero, 75 keV, and 500 keV detection thresholds are denoted as ${Q}_{0}$, ${Q}_{75}$, and ${Q}_{500}$. The energy deposited as light calorimetry under a uniform $\overline{\text{LY}}$ of 220 PE/MeV, 100 PE/MeV, and 35 PE/MeV are denoted as ${L}_{220}$, ${L}_{100}$, and ${L}_{35}$.
Fig. 3 shows the distributions of deposited charge and light from 35 MeV $\nu_{e}$ CC interactions on Ar. In addition to ${E}_{\text{avail}}$ and ${E}_{\text{depo}}$ described in Sec. II, the energy deposited as charge at zero charge detection threshold, ${Q}_{0}$, primarily peaks around 19 MeV (Fig. 3 top plot). Secondary peaks in the ${Q}_{0}$ distribution originate from ${E}_{\text{depo}}$ due to events with knockout hadrons. The secondary peak at 18 MeV in ${Q}_{0}$ corresponds to events with neutron emission, while the broader peak at 12 MeV consists of events with a proton or an $\alpha$ particle. Compared to ${E}_{\text{depo}}$, the Birks model introduces additional broadening in ${Q}_{0}$. This broadening is more pronounced for secondary peaks because knockout hadrons are not MIPs, leading to greater variation in recombination effects as described by the Birks model.
@@
For energy deposited as light, the events are centered around 14 MeV (Fig. 3 middle plot). The secondary peaks are less distinct and merge with the primary peak. This effect is further illustrated in Fig. 4, which shows the energy deposited in light for $\overline{\text{LY}}$ = 220 PE/MeV (${L}_{220}$). Specifically, for events with neutrons, the net bias in deposited energy is --1.8 MeV. The total deposited energy of these events typically fall within the two-standard-deviation band of the primary light peak. Consequently, neutron emission events populate the lower tail of the $\text{L}_{220}$ distribution in Fig. 4 rather than forming a distinct secondary peak, as observed in the charge distribution.
@@
It is important to note that nuclear quenching effects in hadron-nucleus interactions are not explicitly simulated in this study. In particular, a lower scintillation efficiency has been observed in tens of keV nuclear recoils in LXe [44]. For $\alpha$ particles, a smaller light production is observed possibly explained by a biexcitonic quenching process; while for few MeV protons, this was calculated to be negligible [45]. For neutrons in this study, the potential extra light loss could shift neutron emission events further from the primary peak in the light distribution.
The single peak structure of the light distribution is observed across all $\nu_{e}$ samples from 5 to 50 MeV and holds true for all benchmark $\overline{\text{LY}}$. Unlike the charge calorimetry, no systematic energy bias is introduced when varying $\overline{\text{LY}}$. However, energy smearing effects become more pronounced as $\overline{\text{LY}}$ decreases from 220 PE/MeV to 35 PE/MeV. In Fig. 3 middle plot, the standard deviations are 0.56 MeV, 0.59 MeV, and 0.67 MeV for ${L}_{220}$, ${L}_{100}$, and ${L}_{35}$ respectively.
Finally, the event ${E}_{\text{depo}}$ is largely recovered when combining charge and light calorimetry, ${Q}_{0}$ + ${L}_{220}$ (Fig. 3 bottom plot). Similar to charge-only calorimetry, the application of a 75 keV charge detection threshold introduces spectral broadening and an energy bias in the combined calorimetry. Additionally, implementing the modified Box model[46] to simulate recombination effects yields similar results to those obtained using the Birks model.
@@ IV.1 Energy Reconstruction
In this section, we examine the potential for energy reconstruction using charge and light calorimetry in a future LArTPC with an enhanced photon detection system. The ${E}_{\text{avail}}$ for $\nu_{e}$ CC events can be reconstructed by applying a scaling factor to the deposited energy. The scaling factor represents the fraction of ${E}_{\text{avail}}$ that is converted into deposited energy in charge, light, or combined calorimetry in each event. As an example, we consider charge calorimetry with a detection threshold of 75 keV, ${Q}_{75}$. From simulation, we determine a scale factor of 0.54, taken as the mode of the ratio distribution between ${Q}_{75}$ and ${E}_{\text{avail}}$. The reconstructed energy from ${Q}_{75}$ is then given by:
@@
Below we show three analyses (analysis A, B, and C) that go from the most basic to the most difficult scenario regarding the treatment of hadrons. The energy reconstruction capability of the light system is demonstrated through these analyses. We also note that low-energy radioactive backgrounds are not considered in any of the three analyses. Identifying signal-related energy depositions across the detector volume in a high-multiplicity, low-energy environment could present significant challenges if the rate of background activities is very high. Background energy deposits from internal or external radiological sources may be misidentified or mis-clustered as signal deposits. A detailed simulation incorporating both signal and radiological backgrounds is necessary to evaluate the detector’s ability to separate events and assess physics sensitivity, using both charge and light detection. However, such a study is beyond the scope of this paper and warrants a separate investigation. In the analysis presented below, we assume that signal-related energy depositions are efficiently identified throughout the detector volume, with effective charge-light matching.
IV.1.1 Analysis A - Exclude energy deposits from neutron capture
In a realistic LArTPC, there will be inefficiency at tagging neutron captures since they happen much later (more than one drift time) than the initial neutrino interaction. It is straightforward for physics analysis not to include energy deposits from neutron captures. Therefore, for the most basic analysis, we exclude all energy deposits from neutron captures.
The deposited energy distribution after excluding neutron energy deposits is shown in Fig. 5 overlaid with some distributions from Fig. 3 for comparison. For the total deposited energy (Fig. 5 top plot), the previous distinct neutron emission events now cluster close with the other hadron emission events around 26 MeV. After taking into account energy partition based on Birks model, the neutron emission events now peak at around 14.5 MeV in charge calorimetry, as shown in the middle plot of Fig 5. Finally, due to the loss of the capture energy deposits, the neutron emission events cluster at a distinct secondary peak at around 11 MeV in the light calorimetry in contrast to Fig. 4.
Figure 5: Deposited energy distribution for the 35 MeV $\nu_{e}$-Ar CC sample after excluding energy deposits from neutron captures, overlaid with the relevant inclusive energy distribution in Fig. 3. Top: total deposited energy. Middle: energy deposited into charge calorimetry ${Q}_{0}$. Bottom: energy deposited into light calorimetry ${L}_{220}$.